OpenStudy (anonymous):

Solve each equation for 0 = x = 2π : i) 2cos^(2)x + 3cosx – 2 = 0 ii) sinx = sqrt 3cosx (Rewards for correct answer)

OpenStudy (anonymous):

did you mean \[ (1)\quad2cos^2x+3cosx−2=0\\ (2)\quad \sin x=\sqrt{3}\cos x\\ \text{for }0\le x \le2\pi\] ???

OpenStudy (anonymous):

a. (2cosx-1)(cosx+2)=0 cosx=1/2 or cosx=-2 x={π/3, 5π/6} This is correct, i promise...

OpenStudy (zehanz):

If you set p=cos x in the first one, you get: 2p²+3p-2=0. You can solve this, using the quadratic formula. Once you've got you p, remember cos x = p, so solve for x.

OpenStudy (anonymous):

Do you get mines @jahvoan

OpenStudy (anonymous):

i got the answer but what is the corresponding angle for -2 on the unit circle?

OpenStudy (zehanz):

There isn't. cosx=-2 has no solutions, because the x-coordinate of a point on the unit circle (which is the cosine) only has values from -1 to 1...

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

solve equation 2

OpenStudy (zehanz):

You can divide both sides by cos x and remember sinx/cosx = tan x

OpenStudy (anonymous):

but on the right it is sqrt cosx

OpenStudy (zehanz):

No, it's sqrt(3) * cosx

OpenStudy (anonymous):

no look at the problem up top

OpenStudy (zehanz):

It says: sinx = sqrt 3cosx This can only mean:\[ \sin x = \sqrt{3} \cdot \cos x\]So dividing by cos x gives:\[\frac{ \sin x }{\cos x }=\sqrt{3} \Leftrightarrow \tan x=\sqrt{3}\]And this is a "nice" value of tan x. If you interpret it otherwise, you can't solve it...

OpenStudy (anonymous):

so what will be the answer for sqrt 3 on the unit circle?

OpenStudy (zehanz):

Remember that special triangle with sides of a, 2a and a√3?

OpenStudy (anonymous):

is it pi/6

OpenStudy (anonymous):

no its pi/3 right or wrong?

OpenStudy (zehanz):

60 degrees, or pi/3

OpenStudy (anonymous):

thanks mayne

OpenStudy (anonymous):

why is equation 2 not considered to be an identity?

OpenStudy (zehanz):

Identities hold for every x. This one is only true if x = pi/3. Identity: sinx=cos(pi/2-x). No matter what x is, it's always true!